OpenStudy (anonymous):

Help with some math review questions? I don't understand and need someone to explain it to me?

OpenStudy (anonymous):

Solve the following quadratic equations by completing the square: 1. x^2 - 16 = 6x

OpenStudy (anonymous):

2. x^2 + 1= 3x

OpenStudy (anonymous):

3. x^2 = 9 - 7x

OpenStudy (unklerhaukus):

ok. so we have \[x^2 - 16 = 6x\]move all the \(x\) terms to the left hand side

OpenStudy (unklerhaukus):

can you do that ?

OpenStudy (anonymous):

I need help, I'm so sorry, I've totally forgotten all of this

OpenStudy (unklerhaukus):

\[x^2−16=6x\] if we move all the\(x\) terms to the left hand side of the equation \[x^2-6x−16=0\] and move the other terms to the right hand side \[x^2-6x=16\]

OpenStudy (unklerhaukus):

now what could we add to both sides to make the left hand side a perfect square?

OpenStudy (anonymous):

x^2? I'm so confused

OpenStudy (unklerhaukus):

we are looking for a^2

OpenStudy (unklerhaukus):

so that we can use this \[(x-a)^2\\=(x-a)(x-a)\\=x(x-a)-a(x-a)\\=x^2-xa-ax+a^2\\=x^2-2ax+a^2\]in reverse

OpenStudy (unklerhaukus):

can you find \(a\)?